OpenStudy (anonymous):

A student bought 1 box of crayons and 5 reams of paper for 54$ . She bought 5 boxes of crayons and 3 reams of paper for 50$. What is the cost of each box of crayon and each ream of paper

OpenStudy (campbell_st):

you need to solve this simultaneously c = crayons, p = reams of paper so c + 5p = 54 5c + 3p = 50 this can be solved by substitution take the 1st equation and make c the subject.. x = 54 - 5p substitute it into the 2nd equation 5(54 - 5p) + 3p = 50 distribute and simplify 270 - 25p + 3p = 50 - 22p = -220 I'll let you solve for p.... this will be the cost of 1 ream of paper... when you know the cost... substitute into either of the original equations to find the cost of a box of crayons...